LAB PROBLEM 1: MEASURING SPRING CONSTANTS TWO WAYS

Can you help me to solve these questions basically?
Answer #1

When the spring is in unstretched position then restoring force acting on mass (Say m) due to spring will be exactly balanced by the force due to gravity on the mass.

Let its equilibrium position be yo. then using balance of forces,

  F_s+ F_g=0

  -ky_o= mg

where k= spring constant

g= acceleration due to gravity

or y_o= -\frac{mg}{k} ................................(1)

Now let mass be stretches to a new position, say, yn. Then it will have net acceleration in upward direction due to more spring restoring force acting on it. Thus we'll have

F_s-F_g= ma (Using balance of forces and Newton's Laws of motion)

-ky_n-mg= ma(Taking upward direction as positive)

-k(y_n+\frac{mg}{k})= ma

using equation 1 in above,

-k(y_n-y_o)= ma

let yn-yo= y = net displacement from its equilibrium position.

-ky= ma

also acceleration in vertical direction can be written as a= \frac{d^2y}{dt^2}

then we'll have

  m\frac{d^2y}{dt^2}= -ky

\frac{d^2y}{dt^2}+\frac{k}{m}y=0............................(1)

Let the solution of above equation be y= Asin\omega t .....................(2) ( where A= some constant)

then \frac{dy}{dt}= -A\omega cos\omega t

and \frac{d^2y}{dt^2}= -A\omega^2 sin\omega t= -\omega^2 y .........................(3)

Putting equation 2 and 3 in equation 1, we get

-\omega^2 y+\frac{k}{m}y=0

-\omega^2 +\frac{k}{m}=0

we'll get angular frequency of this oscillation as

  \omega= \sqrt{\frac{k}{m}}(ANS)

And if the Time Periof of oscillation is "T" then angular frequency in terms of Time Period is written as

\omega= \frac{2\pi }{T}(ANS)